# How and Why LLMs Use Deprecated APIs in Code Completion? An Empirical Study

### v2

@@ VII Conclusion
In this work, we conducted an evaluation study to investigate the statuses and causes of deprecated API usages in LLM-based code completion. The study results all evaluated LLMs encounter challenges in predicting plausible API usages and face issues with deprecated API usages, influenced by the distinct code context characteristics of the prompts and the characteristics of API deprecations during the evolution of these libraries. We propose two lightweight fixing approaches to mitigate the deprecated API usages and can serve as baselines for future research. We also provide implications for the research directions for the combination of library evolution and LLM-driven code completion and software development.

### v3

@@ VII Conclusion
In this work, we conducted an evaluation study to investigate the statuses and causes of deprecated API usages in LLM-based code completion. The study results all evaluated LLMs encounter challenges in predicting plausible API usages and face issues with deprecated API usages, influenced by the distinct code context characteristics of the prompts and the characteristics of API deprecations during the evolution of these libraries. We propose two lightweight fixing approaches to mitigate the deprecated API usages and can serve as baselines for future research. We also provide implications for the research directions for the combination of library evolution and LLM-driven code completion and software development. We released the code and data of our study at the website [91].
Acknowledgement
This research project is supported by the National Key R&D Program of China (2023YFB4503805), the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), DSO National Laboratories under the AI Singapore Programme (AISG2-GC-2023-008), National Natural Science Foundation of China under Grant No. 62402342, Shanghai Sailing Program (No. 24YF2749500), and the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG96/23).